A student found that a brass cylinder weighs 75.04 g in air and 66.11 g in water (density is 0.998 g/cm³). What is the density of brass?


1
Expert's answer
2021-03-12T05:56:13-0500

The cylinder displaces (74.04g"-" 66.11g)of water.


=7.93g of water.


Volume of Cylinder ="\frac{Mass of displaced water} {Density}"


"=\frac{7.93g}{0.998g/cm^{3}}"


Volume=7.9459cm"^{3}"


Density of Brass ="\frac{Weight of the cylinder} {Volume}"


"=\frac{74.04g}{7.9459cm^{3}} =9.318g/cm^{3}"